OpenStudy (anonymous):

For what numbers x, -2π ≤ x ≤ 2π, does the graph of y = csc(x) have vertical asymptotes? For what numbers x, -2π ≤ x ≤ 2π, does the graph of y = cot(x) have vertical asymptotes?

OpenStudy (jamesj):

By definition csc(x) = 1/sin(x). Hence csc(x) = 1 when sin(x) = 1. So when does sin(x) = 1 in that interval?

myininaya (myininaya):

csc(x) has vertical asymptotes when sin(x)=0 since csc(x)=1/sin(x) cot(x) has vertical asymptotes when sin(x)=0 since cot(x)=cos(x)/sin(x)